Simplify the following expression.
2x - 2y + 5z - 2x - y + 3z

Answer by lenny460(1073)

Simplify the following expression.
2x - 2y + 5z - 2x - y + 3z

Combine like terms:
2x - 2x = 0
-2y - y = -3y
5z + 3z = 8z
Now we have:
-3y + 8z

Answer:
8z - 3y








Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com